OpenStudy (study_buddy99):

simplify the expression (logs) (alg 2) (will medal those who help me understand)

OpenStudy (daniel.ohearn1):

Hi, go ahead

OpenStudy (daniel.ohearn1):

What's the problem?

OpenStudy (study_buddy99):

\[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\]

OpenStudy (daniel.ohearn1):

Would you agree that (a/b) * (c/d) would equal ac/bd?

OpenStudy (study_buddy99):

Idk what that means

OpenStudy (daniel.ohearn1):

a b c and d are letters that represent numbers called variables because we don't what numbers they are

rishavraj (rishavraj):

\[a^x \times a^y = a^{x + y}\] \[\frac{ a^x }{ a^y } = a^{x - y}\]

OpenStudy (daniel.ohearn1):

If I multiply two fractions all I need to do to get the product is multiply across.. 1/2 times 3/4 is an example, what would that be?

OpenStudy (study_buddy99):

3/8

OpenStudy (daniel.ohearn1):

Right so first step, multiply accross..

OpenStudy (study_buddy99):

20e^-20?

OpenStudy (daniel.ohearn1):

rishavraj left a formula that will be useful when multiplying terms with exponents

OpenStudy (study_buddy99):

what? 0.0

OpenStudy (mathmale):

"\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\] is an expression, not an equation, so NO cross mult applies here. Neither can you "solve" for anything. You can only reduce the given expression.

OpenStudy (mathmale):

In your shoes I would combine \[e ^{-4 }~and~e^5 \] thru multiplication; result is just e.

OpenStudy (mathmale):

So, reduce the given expression as much as possible.

OpenStudy (study_buddy99):

what do you do with the 10 and the 2?

OpenStudy (mathmale):

In the expression \[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\] you have 10 in the numerator and 5 in the denom. Reduce this fraction.

OpenStudy (mathmale):

Can't do anythign with the 2 in the numerator and the 3 in the den. other than keep them as is.

OpenStudy (study_buddy99):

that still makes no sense

OpenStudy (daniel.ohearn1):

2 and 10 you can multiply across to get 20 just as you can with the numberical portion of the denominator but when you e^-4 and you are multiplying by e^5 instead of multiply the exponents we add them say it was 2^5 times 2^2 that will not equal 2^10 that will equal 2^7

OpenStudy (mathmale):

Re: "that still makes no sense:" Rather than say that, would you please try to phrase questions whose answers might make things less confusing to you. "that still makes no sense" leaves the problem dumped into the other person's lap.

OpenStudy (daniel.ohearn1):

So multiply through numerically and add through left to right with the terms of e

OpenStudy (mathmale):

10/5 = 2, right? so the 10 and 5 disappear and are replaced by a 2 in the numerator. 2/3 stays as it is (this fraction can't be reduced).

OpenStudy (study_buddy99):

you said we aren't cross multiplying

OpenStudy (mathmale):

so, where are we now? What do y ou need to know to complete the simplification of the expression \[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }~??\]

OpenStudy (daniel.ohearn1):

Know you are multiplying through right? a b / cd = a / c times b / d

OpenStudy (mathmale):

Cross multiplication? first, I ask you how cross mult. might help and why you think it applies here. Cross mult. has its place when you are working with equations involving fractions. You don't have an equation here. You have an expression which involves fractions.

OpenStudy (mathmale):

therefore, no cross multiplication. Instead, reduce ()which involves division, not mult.).

OpenStudy (mathmale):

\[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\]

OpenStudy (daniel.ohearn1):

In this situation we have a over b times c /d in that each term is unique so you see the process we must take?

OpenStudy (mathmale):

1. Divide that 5 into that 10. 2. Combine e^(-4) and e^2 into ONE exponential, to appear in the numerator. 3. Combine e^5 and e^(-3) into ONE exponenetial, to appear in the numerator. 4. Combine the 2 expressions in the numerator that involve exponentiation.

OpenStudy (study_buddy99):

2e^7 is what I got from 10e^5/5e^2

OpenStudy (daniel.ohearn1):

Do you agree we'd have this if we multiply through? \[(2e ^{-4})(10e ^{5}) / (5e ^{2})(3e ^{-3})\]

OpenStudy (daniel.ohearn1):

Multiply across left to right

OpenStudy (mathmale):

Can't reduce the 2/3. Start your "answer" by typing \[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\rightarrow \frac{ 2 }{ 3 }*\frac{ 2e^1 }{ e ^{-1} }\]

OpenStudy (mathmale):

As daniel suggested, I am multiplying from left to right. \[e ^{-4}e^5=e^1=e\]

OpenStudy (daniel.ohearn1):

So we'll have 20e^( ) / 15e^ ( ) Can you fill in the blank knowing a^x times a^y = a^ x + y for any number value a

OpenStudy (study_buddy99):

I don't know where you're getting e^-1

OpenStudy (daniel.ohearn1):

e^2 times e ^-3

OpenStudy (study_buddy99):

then where are you getting 10/5

OpenStudy (daniel.ohearn1):

On the bottom we have 5e^2 (3e^-3) We multiply the numbers and add the exponents

OpenStudy (mathmale):

Please make a table of Rules of Exponentiation:\[a^b a^c = a ^{b+c}\]...and so on.

OpenStudy (mathmale):

\[a^b a^c=a ^{b+c}\]

OpenStudy (mathmale):

...and so on. You'll need to know these rules.

OpenStudy (mathmale):

"Where did y ou get the 10 and 5?" Please refer back to the original problem statement, the expression \[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\]

OpenStudy (mathmale):

The 10 is in the numerator and the 5 in the denom.

OpenStudy (study_buddy99):

right, so that makes 2

OpenStudy (mathmale):

Yes, it does. 2/3 can't be reduced, so leave that as is. Mult. 2 and (2/3), you get (4/3). OK with that?

OpenStudy (mathmale):

Then you'll need to apply rules of exponentiation to simplify the rest of this expression, which consists of exponential quantities.

OpenStudy (study_buddy99):

where are you getting the other 2 from?

OpenStudy (daniel.ohearn1):

(4/3) is the same as (20/15) which is what you get when you multiply left to right

OpenStudy (mathmale):

Please refer to the original expression:\[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\]

OpenStudy (mathmale):

\[\frac{ 2e ^{-4}}{ 5e ^{2} }*\frac{ 10e ^{5} }{ 3e ^{-3} }\]

OpenStudy (mathmale):

That "2" comes from dividing 10 by 5. We've discussed that several times. The (2/3) comes from the 2 in the numerator and the 3 in the denom.

OpenStudy (mathmale):

So, again, you have (4/3) times a fraction involving powers of e. Combine all those powers of e into one exponential expression.

OpenStudy (study_buddy99):

okay so I got 4e/3e^-1?

OpenStudy (mathmale):

Yes. But that could be simplified to look much better: \[\frac{ 4 }{ 3 }\frac{ e }{ e ^{-1} }=?\]

OpenStudy (study_buddy99):

4/3 *1/e?

OpenStudy (mathmale):

simplify that. this is the last step. Please, if you've written down those rules of exponents as I've suggested, use the appropriate one to simplify \[\frac{ e }{ e ^{-1} }.\]

OpenStudy (mathmale):

Your most result result is almost, but not quite, right. Reduce:\[\frac{ e }{ e ^{-1} }.\]

OpenStudy (study_buddy99):

what rules are you talking about?

OpenStudy (mathmale):

Study_Buddy: I've mentioned rules of exponentiation at least twice in our previous exchanges, recommending that you look them up and copy them down for later reference. I'm a bit concerned because you apparently haven't seen or understood what I'm asking you do do, which has apparently led to your asking questions that have already been answered. I gave you the rule\[e^a e^b=e ^{a+b}\]

OpenStudy (mathmale):

Several others follow:\[e^a/e^b=e ^{a-b}\]

OpenStudy (mathmale):

\[(e^a)^b=e ^{ab}\]

OpenStudy (mathmale):

Please, do an internet search, or look in the index of your algebra textbook, for "exponents, rules of" ... and copy down these rules, study them and review them from time to time.

OpenStudy (study_buddy99):

I don't understand the rules. I don't understand what they represent. I don't understand when they use them.

OpenStudy (mathmale):

think: 1) multiplication of exponential functions 2) division of exponential functions 3) powers of exponential functions do you have a textbook available? If so, would you please look up "exponents, rules of?" Otherwise do the same search on the Internet. If you want my help in deciding which rule to apply in a given case, find some homework problems involving exponential functions and share them with me. Before doing so, compare the problems with the several rules of exponents that I gave you and make a preliminary decision regarding which one to use for which case. It is essential that you learn to do this (practice selecting the appropriate rule to apply to a given homework problem).

OpenStudy (mathmale):

Is there any way in which you could find a live tutor with whom to interact when you feel stuck like this? I have the feeling you'd understand me a bit better were we meeting face to face, not through the Internet.

OpenStudy (mathmale):

I will be back on OpenStudy tomorrow morning, in case you decide to share any more questions with me. I need to get off the 'Net now.

OpenStudy (study_buddy99):

yeah, I'll see someone tomorrow. Thank you for being so persistent with me, I've never seen these rules in my life.